OpenStudy (anonymous):

Which fractions are equivalent to -4/9 ? Choose all answers that are correct. A. -2/5 B. -8/18 C. 16/-36 D. 20/45

OpenStudy (anonymous):

@mathmale i need help with a masth question please

OpenStudy (susanna):

\[-\frac{ 4 }{ 9 }\]

OpenStudy (susanna):

It's not A

OpenStudy (anonymous):

thxs how u get that

OpenStudy (anonymous):

sorry i looked a it wrong

OpenStudy (susanna):

\[-\frac{ 8 }{ 18 }=-\frac{ 4 }{ 9 }\] If you divide both the denominator and numerator by 2, thus simplifying it.

OpenStudy (anonymous):

ohh thxs

OpenStudy (susanna):

u'r welcome

OpenStudy (anonymous):

umm its choose MORE than one

OpenStudy (susanna):

ok

OpenStudy (susanna):

\[-\frac{ 16 }{ 36 }=-\frac{ 4 }{ 9 }\]

OpenStudy (anonymous):

ok thxs

OpenStudy (susanna):

A and D are wrong

OpenStudy (susanna):

no prob

OpenStudy (anonymous):

got another question

OpenStudy (susanna):

sure